OpenStudy (anonymous):

Find all the zeros (please show work). x^4 - 8x^2 - 9 = 0

OpenStudy (anonymous):

put x^2=y y^2-8y-9=0 1*-9=-9 9-1=8 9*-1=-9 write -8y=-9y+1y make factors and put y=x^2 again if possible make factors.

OpenStudy (anonymous):

Put x^2 = X -> X must be positive Solve X^2 - 8X - 9 = 0. Since a - b + c = 0, use the TIP 2 to get the 2 real roots: X1 = -1 (rejected), and X = (-c/a) = 9 x^2 = X = 9-> x = -3 (rejected), and x = 3.